The probability density function for random variable Wis given as follows: 120 w>0 20 Let x be the 100pth percentile of W and y be the 100(1-p)th percentile of W, where o<p〈1. Express y as a function of x. ln(1-e- x/20 ) 20 -x/20 In 1-e 20 Cy- -20 ln 1-ex20 y20 n e-/20
Answer #1

P(W < x) = p

1 - e^(-x/20) = p

P(W > y) = p

e^(-y/20) = p

hence

1 - e^(-x/20) = e^(-y/20)

y = -20 ln (1 - e^(-x/20))

option C) is correct

Please rate